# Nicolas Crampé1, Rafael I. Nepomechie2, Luc Vinet3,4,5

### v1

@@
We show that, after suitably adjusting a uniform transverse magnetic field, the generic inhomogeneous open XX spin chain has a two-fold degeneracy, and an exact $su(2)$ symmetry whose “inhomogeneous” generators depend on coefficients that can be explicitly computed for models associated with discrete orthogonal polynomials.
@@ 1 Introduction
where $J_{n}$ and $B_{n}$ are real parameters. This Hamiltonian has the evident $U(1)$ symmetry
@@
where $\mathbb{S}^{x}$ and $\mathbb{S}^{y}$ are corresponding “inhomogeneous” generators of the form
@@
Generators of $su(2)$ that depend on model parameters and are spatially dependent appeared recently in [6]. However, the system considered there is homogeneous, and the symmetry is only up to boundary terms.
@@ 4 Discussion
We have seen that, after suitably adjusting a uniform transverse magnetic field, the generic inhomogeneous open XX spin chain has a two-fold degeneracy, and an enhanced $su(2)$ symmetry (1.3). The “inhomogeneous” symmetry generators (1.4) depend on coefficients $a_{n}$ (2.15) that can be computed explicitly for models associated with discrete orthogonal polynomials. The expressions for the generators (1.4), with their non-local “tails” of $\sigma^{z}$ operators, are reminiscent of coproducts of quantum-group generators [9]. The symmetry uncovered here may provide further insight into protected phases in inhomogeneous spin chains [10]. It would be very interesting if this type of symmetry could be extended to XXZ-type chains. It would also be interesting to find higher-rank algebras with such “inhomogeneous” generators, and corresponding physical models with such symmetries.

### v2

@@
We show that, after suitably adjusting a uniform transverse magnetic field, the generic inhomogeneous open XX spin chain has a two-fold degeneracy, and an exact $su(2)$ symmetry whose “inhomogeneous” nonlocal generators depend on coefficients that can be explicitly computed for models associated with discrete orthogonal polynomials.
@@ 1 Introduction
where $J_{n}\neq 0$ and $B_{n}$ are real parameters. This Hamiltonian has the evident $U(1)$ symmetry
@@
where $\mathbb{S}^{x}$ and $\mathbb{S}^{y}$ are corresponding “inhomogeneous” nonlocal generators of the form
@@
Generators of $su(2)$ that depend on model parameters and are spatially dependent appeared recently in [6]. However, the system considered there is homogeneous, and the symmetry is only up to boundary terms. The generators (1.4) resemble inhomogeneous versions of so-called semilocal charges considered in e.g. [7, 8].
@@ 4 Discussion
We have seen that, after suitably adjusting a uniform transverse magnetic field, the generic inhomogeneous open XX spin chain has a two-fold degeneracy, and an enhanced $su(2)$ symmetry (1.3). The “inhomogeneous” symmetry generators (1.4) depend on coefficients $a_{n}$ (2.15) that can be computed explicitly for models associated with discrete orthogonal polynomials. The expressions for the generators (1.4), with their non-local “tails” of $\sigma^{z}$ operators, are reminiscent of coproducts of quantum-group generators [11]. The symmetry uncovered here may provide further insight into quench dynamics [7, 8] and protected phases [12] in inhomogeneous spin chains. It would be very interesting if this type of symmetry could be extended to XXZ-type chains. It would also be interesting to find higher-rank algebras with such “inhomogeneous” generators, and corresponding physical models with such symmetries.